Grade 12th passPhysical Chemistry

A mixture is prepared by mixing 10gm H2SO4 and 40gm SO3 find mole fraction of H2SO4?

Prasoon
9 Years agoGrade 12th pass

1 Answer

Shreya
ApprovedApproved Tutor Answer9 Years ago
Mole fraction=mole of h2so4/total moles H2SO4 MOLES .=10/98 =0.102So3 moles=40/80=0.5Mole fraction =0.102/0.602=0.1685